A polarizer and analyzer are oriented so that the amount of transmitted light is maximum. 

Through what angle should either be turned so that the intensity of transmitted light is 

reduced to (a) 0.75 and (b) 0.25 times the maximum intensity?


1
Expert's answer
2022-02-02T17:34:40-0500

"I=I_0\cos^2\varphi,"

"\cos\varphi=\sqrt{\frac I{I_0}},"

a)

"\varphi_1=\arccos\sqrt{0.75}=30°,"

b)

"\varphi_2=\arccos\sqrt{0.25}=60°."